Calculate magnitudes of the components of a vector V if the magnitude of V is 10m and the angle between the vector an the X axis is 45 ?


1
Expert's answer
2021-10-08T11:44:11-0400

"V_y=V\sin\alpha=7.1~m,"

"V_x=V\cos\alpha=7.1~m."